OpenStudy (slayer8001):

What is the solution set of the following equation? -8x + x + 15 = -7x + 12

OpenStudy (anonymous):

-8x+x+15=−7x+12 Step 1: Simplify both sides of the equation. −8x+x+15=−7x+12 Simplify: (Show steps) −7x+15=−7x+12 Step 2: Add 7x to both sides. −7x+15+7x=−7x+12+7x 15=12 Step 3: Subtract 15 from both sides. 15−15=12−15 0=−3 Answer: There are no solutions.

OpenStudy (slayer8001):

ok one more @aloud Solve for n. -53 + n = -28

OpenStudy (anonymous):

−53+n=−28 Step 1: Simplify both sides of the equation. n−53=−28 Step 2: Add 53 to both sides. n−53+53=−28+53 n=25 Answer: n=25

OpenStudy (slayer8001):

oops my bad this one is the last one

OpenStudy (anonymous):

ok

OpenStudy (slayer8001):

Find two consecutive integers whose sum is 67. Which of the following equations could be used to solve the problem? 2x=67 2x+1=67 2x+2=67 x^2+1=67

OpenStudy (anonymous):

im not sure but i think its x^2+1=67